OpenStudy (anonymous):

Rachel deposited $5,960.32 into a savings account with an interest rate of 4.2% compounded twice a year. About how long will it take for the account to be worth $9,000?

OpenStudy (anonymous):

The formula is: N = ln(fv/pv)/ln(1+i) i = 4.2 / (2 x 100) = 0.021 1+i = 1.021 fv /pv = 9000/ 5960.32 = 1.51 N = ln(1.51)/ln(1.021) = 0.4121 / 0.0208 =19.829 6 month periods = 9.915 years